OpenStudy (96mertcank96):

The probability that the cubs win their first game is 1/3 .the probability that the cubs win their second game is 3/7 . what is the probability that the cubs win both games?

OpenStudy (ivancsc1996):

The multiplication of both probabilities:\[\frac{ 1 }{ 3 }\frac{ 3 }{ 7 }=\frac{ 1 }{ 7 }\]

OpenStudy (96mertcank96):

i dont have an answer choice that has 1/7

OpenStudy (anonymous):

You are supposed to add them I believe @96mertcank96 , so that would be 16/21

OpenStudy (96mertcank96):

thank you again!

OpenStudy (anonymous):

Your welcome :) @96mertcank96